To estimate the size of the bear population, conservationists, captured, tagged and released 50 bears. One year later a random sample of the 100 bears included only 2 tagged bears. What is the conversationist's estiamte of the size of the bear population?
I love ratio questions but I am stumped on this one. I only have the follwing:
2/100=

Found 2 solutions by user_dude2008, solver91311:
Answer by user_dude2008(1862)

tagged/total (second year) = tagged/total (first year)
2/100=50/x
2x=50*100
2x=5000
x=5000/2
x=2500 <------ Answer

Answer by solver91311(24713)



2 percent of the 100 bears surveyed were found to be tagged, so you can make the assumption that 2 percent of the total population was tagged. Since 50 bears were tagged, 50 equals .02 times the total population.

Solve:




John